OpenStudy (anonymous):

solve for cot 2B = 4 tan B

OpenStudy (earthcitizen):

solve for what ?

OpenStudy (anonymous):

find B given 0<B<360

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=cot+2B+%3D+4+tan+B+

jhonyy9 (jhonyy9):

us cotB=1/tanB

OpenStudy (earthcitizen):

\[\tan(2\theta)= 2\tan \theta/1-\tan ^{2}\theta\]

OpenStudy (earthcitizen):

cot2b=1/tan2b

OpenStudy (earthcitizen):

cot2b=cos2b/sin2b

OpenStudy (earthcitizen):

\[\cot2b=(1-\tan ^{2}b)/(2tanb)\]

OpenStudy (anonymous):